# Guardians and Offenders: A Survey on Harmful Content Generation and Safety Mitigation

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
@@ 2.1. Categorization of harmful content in LLM
We systematically categorize the harmful content generated by LLM into four types according to their impact mechanisms, as shown in Table 1. Toxic content encompasses both explicit abuse—such as profanity or slurs—and implicit hostility, which relies on rhetorical devices (e.g., euphemism, sarcasm, metaphor) and shared commonsense or social‐norm knowledge to convey harm (Zhang et al., 2023) (284, ). Harassment refers to directed interpersonal attacks—bullying, doxxing threats, or group‐based verbal aggression—aimed at intimidating or silencing individuals or collectives (Vidgen et al., 2019) (vidgen2019much, ). Offensive language spans explicit hateful remarks that are overt and relatively easy to detect, as well as implicit offenses whose harm emerges only in specific contexts or against particular targets (Das et al., 2024) (das2024offensivelang, ). Finally, bias content targets individuals or groups based on protected or socially salient attributes—such as race, gender, religion, ethnicity, or sexual orientation—and perpetuates systemic inequalities (Albladi et al., 2025) (xu2024hatespeech, ). Recognizing these categories is crucial for developing targeted detection and mitigation strategies within LLM.
@@ 2.2. Terms and definitions
Text generation—known as natural language generation—involves producing coherent, human-like text by predicting and decoding sequences of tokens; effective decoding strategies elevate LLMs from mere next-token predictors to versatile generative engines (Becker et al., 2024 (becker2024text, ); Li et al., 2022 (li2024pre, ); Nabi et al., 2024 (LLM-text-generation, )). Text classification is the task of assigning predefined labels—such as sentiment or topic—to an input text, forming the backbone of many downstream filtering and analysis pipelines (Wang et al., 2023 (sun2023text, ); Min et al., 2023 (dong2022survey, )). Content moderation refers to the systematic review and filtering of user-generated content to enforce a platform’s policies and community standards, preventing abuse and facilitating cooperation (Industry source, 2023 (tspa2025content, ); Grimmelmann, 2015 (grimmelmann2015virtues, )). Finally, Jailbreak attacks exploit carefully crafted prompts to bypass an LLM’s built-in safety guardrails, coaxing prohibited or harmful outputs and posing a serious security challenge for both research and commercial deployments (Zhou et al., 2024; Yu et al., 2024 (Zhou2024-uh, ; Yu2024-pm, ); Chen et al., 2023 (Chen2023-eq, )). This layered framework enables a comprehensive understanding and systematic addressing of LLM safety challenges, from the fundamental generation mechanisms to the practical implementation of safety measures, while acknowledging both defensive capabilities and potential vulnerabilities. In this paper, we define Prevention as the tool or process to prevent the generation of the toxic content, online harassment, offensive language and biased content, which are defined in Table 1.

### v2

@@ 2.1. Categorization of harmful content in LLM
We systematically categorize the harmful content generated by LLM into four types according to their impact mechanisms, as shown in Table 1. Toxic content encompasses both explicit abuse—such as profanity or slurs—and implicit hostility, which relies on rhetorical devices (e.g., euphemism, sarcasm, metaphor) and shared commonsense or social‐norm knowledge to convey harm  (284, ). Harassment refers to directed interpersonal attacks—bullying, doxxing threats, or group‐based verbal aggression—aimed at intimidating or silencing individuals or collectives (vidgen2019much, ). Offensive language spans explicit hateful remarks that are overt and relatively easy to detect, as well as implicit offenses whose harm emerges only in specific contexts or against particular targets (das2024offensivelang, ). Finally, bias content targets individuals or groups based on protected or socially salient attributes—such as race, gender, religion, ethnicity, or sexual orientation—and perpetuates systemic inequalities (xu2024hatespeech, ). Recognizing these categories is crucial for developing targeted detection and mitigation strategies within LLM.
@@ 2.2. Terms and definitions
Text generation—known as natural language generation—involves producing coherent, human-like text by predicting and decoding sequences of tokens; effective decoding strategies elevate LLMs from mere next-token predictors to versatile generative engines (becker2024text, ; li2024pre, ; LLM-text-generation, ). Text classification is the task of assigning predefined labels—such as sentiment or topic—to an input text, forming the backbone of many downstream filtering and analysis pipelines  (sun2023text, ; dong2022survey, ). Content moderation refers to the systematic review and filtering of user-generated content to enforce a platform’s policies and community standards, preventing abuse and facilitating cooperation  (tspa2025content, ; grimmelmann2015virtues, ). Finally, Jailbreak attacks exploit carefully crafted prompts to bypass an LLM’s built-in safety guardrails, coaxing prohibited or harmful outputs and posing a serious security challenge for both research and commercial deployments (Zhou2024-uh, ; Yu2024-pm, ; Chen2023-eq, ). This layered framework enables a comprehensive understanding and systematic addressing of LLM safety challenges, from the fundamental generation mechanisms to the practical implementation of safety measures, while acknowledging both defensive capabilities and potential vulnerabilities. In this paper, we define Prevention as the tool or process to prevent the generation of the toxic content, online harassment, offensive language and biased content, which are defined in Table 1.